OpenStudy (lgbasallote):

LGBARIDDLE If you take away nine from six, then take away ten from nine, and lastly, take away fifty from forty, then you get six. How is this possible?

OpenStudy (lgbasallote):

this one is actually harder

OpenStudy (anonymous):

akitav doesnt it give this: 6 - 9 + 10 + 40 - 50 whjen you expand? it doesnt give you 6 though O_O

OpenStudy (anonymous):

Yes, i missed out the second step, which is why i took it off.

OpenStudy (lgbasallote):

that's where the "riddle" part of "lgbariddle" comes in

OpenStudy (anonymous):

sorry didnt see you took it off man :/

OpenStudy (anonymous):

Math wouldn't be fun if there weren't any wrong answers, then it'd just be philosophy. @MarcLeclair No worries, life is good :)

OpenStudy (lgbasallote):

math is never fun

OpenStudy (konradzuse):

I got it. 40-50 = -10 + 10 = 0 9-9 = 0 + 6

OpenStudy (konradzuse):

goodnight iggy.

OpenStudy (anonymous):

Weird I love cal 1

OpenStudy (lgbasallote):

this is more on out-of-the-box-thinking rather than mathematics

OpenStudy (anonymous):

@lgbasallote Oh wise green one!

ganeshie8 (ganeshie8):

well then you cant take 9 from 6, or any of the rest of the things. so ur 6 will be there as it is

OpenStudy (lgbasallote):

nope. think more out-of-the-box

OpenStudy (anonymous):

I hate yo uthis is haunting me right now

OpenStudy (lgbasallote):

that's what i do

OpenStudy (lgbasallote):

then you feel very bad for missing a simple solution to these riddles. i love those feelings.

OpenStudy (anonymous):

any hint ? :/

OpenStudy (anonymous):

I hate this, I can't figure it out :p. I was thinking 6 groups of things, but I'm probably wrong.

OpenStudy (mimi_x3):

are we allowed to rephrase the question?

OpenStudy (lgbasallote):

rephrase how?

OpenStudy (mimi_x3):

instead of taking away nine from six; we take away six from nine lol my teacher taught me this for combinations and permutations..dont know if its relevant sound stupid

OpenStudy (lgbasallote):

nope you can't

OpenStudy (lgbasallote):

you have to follow directions

OpenStudy (shubhamsrg):

got it!! lol..1 min i shall post..

OpenStudy (lgbasallote):

sure

OpenStudy (shubhamsrg):

SIX – IX = S IX – X = I XL – L = X ------->SIX :D

OpenStudy (shubhamsrg):

roman numerals eh!! nice..

OpenStudy (lgbasallote):

very well young pupil